110 likes | 221 Views
DT Infrastructure: LV, HV, Gas. CMS week 28 June 2011 A. Benvenuti (INFN Bologna). A Few Problems Since May Workshop . 5 Work Packages for LV related problems 6 WP for HV fixes Fan Tray Operation test (with Silverio Dos Santos) => It does not !. LV Problems.
E N D
DT Infrastructure: LV, HV, Gas CMS week 28 June 2011 A. Benvenuti (INFN Bologna)
A Few Problems Since May Workshop • 5 Work Packages for LV related problems • 6 WP for HV fixes • Fan Tray Operation test (with Silverio Dos Santos) => It does not !
LV Problems • In reality only two problems in YB-1: • - A3100 Connector temperature • - A3050 (MB1 MB2 S12) repeated power off due to high temperature interlock • One UXC access was needed because the A3050 had lost communication due to the brutal switch off following the power cut. Was fixed by extracting and reinserting the module • Vcon error on YB+2 MB3S12 only a modest temperature increase (but not the first time)
YB-1 MB1, MB2 S12 LV Supply °C MC OFF (May 24 @ 22:20) • The A3050 switched OFF the MC on a high temperature interlock on May 24 ( 64°C), June 15(66°C) and June 25 (72°C) • The first two instances were recovered quickly • On June 25 , switching ON the MB1 MC caused again the interlock. • The rest of the fill was taken with MB1 S12 MC off • After the beams were dumped, ~13:20, the MB1 went ON, but we decided to replace the A3050 • Looking at the error messages, I discovered two additional cases on February 10 and March 10 • What should we do with this module? We can send it to CAEN but I am afraid they will not be able to reproduce the problem. June 7, Temperature jumped to 37 °C °C MC OFF (June25 @ 07:12) MC OFF (June15 @ 22:16)
Sensors Mounted on A3100 side Sensors mounted on cable connectors at SC side YB-1 • Alarms started on June 24 ~13:00 as the temperature measured with V307 went above 27.7 °C • The temperature stabilized at 29.1 °C . • In UXC It was measured at 41.5 °C • Disconnecting and reconnecting the cable brought V307 down to 25.8 °C • The same incident occurred on May 20 • The A3100 is in the same crate where we substituted the A3050 for MB1MB2 S12
LV Modules Temperatures • The available cooling power is marginal for the crates with A3100 • Check heat exchangers during next Technical Stop • After the insertion trick the A3100 temperature dropped to 39 °C
HV Problems 3 DT stuck/tripped due to HV Spikes on wires while at step 1 YB-1 MB2S02, switched OFF while in STANDBY 16 Channels lost A876 Replaced L2 16 Channels lost 16 Channels lost 5 modules that failed in the detector but were OK in B904 were inserted in easy positions. OK so far
HV Problems In Summary: 12 Chambers: 4MB1, 1MB2, 2MB3, 5MB4 6 in YB0, 2 in YB±1, 1in YB±2 If the detector opens for the Pixels, priority to YB-2, we can recover 1 or 2 MB4
Planned Activities During Next TS Done Improving the TJB grounding consists of a few steps: - Substitute plastic cover with aluminum plate - add pigtail ground connection to TJB - close box with aluminum tape - connect ground shield on the cable with box - wrap everything with aluminum foil… • Reduces synchronous noise (it can’t hurt) • It is difficult to intervene on the TJB (if needed) • 10 TJBs to be done • Make the procedure simpler
Modified A3050s (Matteo) • 4 Modified A3050 are in B904 • Daniele will test them this Friday • We will replace modules in YB0 sectors that already gave problems: • MB1S01, MB4S02, MB2S10, MB3S04
Gas System YB0 H20 content (ppm) from March 08 to May 31 • Stable situation (no changes) • H2O content increased 30% (60% in peak) since March 8 • O2 content in input lines varies from 82 to 106 ppm and from 89 to 103! In output • Position of O2 spikes very stable • N2 content measured in USG on common return line but data still not summarized H20 Content in peak from March 08 to June 14 • It is summer, vacation time • Decided with Roberto to consolidate: • - O2 analysis cells calibration • - H2O station calibration (YB0) • - N2 last measured • In September review some issues: Online gas chromatograph, H20 control (~400 ppm), O2 content (100 - 200 ppm)